Two steel balls A and B of mass 10kg and 10g roll towards each other with 5m/s and 1m/s resp. on a smooth floor. After a perfectly elastic collision with what speed B moves?

Grade:12th pass

1 Answers

Arun
4 years ago
Mass of ball A = 10 kg
Velocity = V1 = 5 m/s
Mass of ball B = 10 g = 10/1000 = 0.01 kg
Velocity = V2 = 1 m/s
Momentum of ball A before collision = (10) (5) = 50 kg m/s
Momentum of ball B before collision = (0.01) (1) = 0.01 kg m/s
Total momentum before collision = 50 + 0.01 = 50.01 kg m/s
Since momentum is conserved, So total momentum after collision must be equal to 50.01 kg m/s
Now,
Total mass = Mass of ball A = Mass of ball B
Total mass = 10 + 0.01 = 10.01 kg
For velocity,
p = m v
v = p/m
v = 50.01/10.01
v = 4.99 m/s or 5 m/s